OpenStudy (anonymous):

Help ? Last year, Yolanda purchased 60 shares of Stock A at $175 per share, 110 shares of Stock B at $100 per share, and a $1250 bond with a 4.75% coupon for $1400. Yolanda sold both stocks today. Stock A is worth $185 per share and Stock B has a value of $104 per share. Assuming neither stock paid a dividend, which investment has the lowest rate of return? Stock A Stock B Bond Stock A and Stock B

OpenStudy (amistre64):

the only money you got from the stock is the money from selling it

OpenStudy (amistre64):

A: 60(185-175) B: 110(104-100) just have to determine the bond

OpenStudy (amistre64):

not too sure what to do about the bond .... 1400 - 1200(1.0475) maybe?

OpenStudy (amistre64):

difference/initial investment is rate of return (185-175)/175 = .057 (104-100)/100 = .040 (1200(1.0475)-1400)/1400 .... this ones iffy still, but i dont think youve recouped the initial 1400 yet

OpenStudy (amistre64):

either that or .0475 is the rate of return for the bond ....

OpenStudy (anonymous):

Stock B ?

OpenStudy (amistre64):

i cant verify for correctness, but yes, stock B seems to have had the lowest rate of return to me. 1200(.0475) = 57 earned on the bond 57/1400 = .0407; which is still a little better on the bond investment ...

OpenStudy (anonymous):

ok , thank you ! :D

OpenStudy (amistre64):

youre welcome